OpenStudy (anonymous):

If f(x) = 3x + 7 and g(x) = 5x – 2 then (f+g)(x) = 8x+5 TRUE OR FALSE

jimthompson5910 (jim_thompson5910):

(f+g)(x) is the same as f(x) + g(x)

jimthompson5910 (jim_thompson5910):

so all you're doing is adding the two functions f(x) and g(x) which really the same as adding 3x+7 to 5x-2

OpenStudy (anonymous):

Okay, so then what?

jimthompson5910 (jim_thompson5910):

add the two expressions and tell me what you get

OpenStudy (anonymous):

okay 8x -9

OpenStudy (zzr0ck3r):

f(x) + g(x) = (3x + 7) + (8x + 5) can you simplify this?

OpenStudy (anonymous):

i think so

jimthompson5910 (jim_thompson5910):

to add 3x+7 to 5x-2, you just combine like terms 3x+5x = ?? 7 + (-2) = ??

OpenStudy (anonymous):

8x + 5

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so them claiming that If f(x) = 3x + 7 and g(x) = 5x – 2 then (f+g)(x) = 8x+5 is a true statement

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome